# Guided Policy Optimization under Partial Observability

### v1

@@ 1 Introduction
Many real-world tasks can be formulated as sequential decision-making problems where agents take actions in an environment to achieve specific goals over time (puterman2014markov, ). Reinforcement Learning (RL) has emerged as a powerful tool for solving such tasks, leveraging trial-and-error learning to optimize long-term rewards (Sutton1998, ). Despite its success, RL encounters significant hurdles in complex and partially observable environments, where agents often operate with limited or noisy information (madani1999undecidability, ). However, during training, we often have access to extra information that could significantly enhance learning efficiency and performance (doi:10.1126/scirobotics.abc5986, ; pmlr-v164-chen22a, ). For instance, in robotics, while real-world sensor data may be noisy or incomplete, simulation environments typically provide full state observability.
@@
Despite the potential of such privileged information, effectively leveraging it in practice remains a major challenge. One popular strategy to utilize this information is through methods like Imitation Learning (IL) (hussein2017imitation, ), Teacher-Student Learning (TSL), or policy distillation (pmlr-v89-czarnecki19a, ). In these approaches, a teacher, equipped with privileged information, provides supervision to guide the student’s learning process. However, this strategy introduces its own set of challenges: a teacher with privileged information may impose an unrealistically high-performance standard, making it difficult for the student to effectively imitate. This issue, known as the “impossibly good” teacher (Walsman2023ImpossiblyGE, ) or imitation gap (10.5555/3540261.3541724, ), can hinder learning and degrade performance. To address this, previous work has sought to integrate environmental rewards into the learning process of the student. One approach is to combine RL with IL (10.5555/3540261.3541724, ; pmlr-v202-shenfeld23a, ; pmlr-v205-nguyen23a, ), switching to RL-based training when the teacher becomes inimitable. Another approach modifies environmental rewards based on the teacher through policy distillation (pmlr-v89-czarnecki19a, ; Walsman2023ImpossiblyGE, ). However, such methods diminish the utility of privileged information, often resulting in inefficient use of the teacher’s knowledge.
@@
To better exploit available information, we propose training a “possibly good" teacher. Inspired by Guided Policy Search (GPS) (pmlr-v28-levine13, ; NIPS2016_a00e5eb0, ), we introduce Guided Policy Optimization (GPO), a novel framework that trains both the teacher and the student simultaneously while ensuring that the teacher’s policy remains aligned with that of the student. The key insight behind GPO is that by leveraging privileged information during training, the teacher can be trained more effectively while ensuring that its performance is “possibly good," thus facilitating easier imitation by the student. Theoretically, we show that the student can achieve optimality similar to direct RL training, mitigating the suboptimality and imitation gaps that often arise from purely teacher-based supervision.
We empirically validate our algorithm across various tasks. In tasks where traditional guidance methods fail to produce optimal policies, our approach proves highly effective. We further validate our algorithm on challenging continuous control tasks in partially observable, noisy environments within the Brax (brax2021github, ) domain, where GPO outperforms baseline methods. Additionally, in memory-based tasks from the POPGym (morad2023popgym, ) benchmark, GPO shows significant improvements, underscoring its ability to exploit extra information and deliver robust performance across diverse domains.
@@ 2 Background
We consider Partially Observable Markov Decision Process (POMDP) (KAELBLING199899, ), which is characterized by the tuple $\langle\mathcal{S},\mathcal{A},r,\mathcal{P},\mathcal{O},\gamma\rangle$. $\mathcal{S}$ represents the set of states, $\mathcal{A}$ the set of actions, $r$ the reward function, $\mathcal{P}$ the transition probability function, $\mathcal{O}$ the partial observation function and $\gamma$ the discount factor. At each time step $t$, the agent receives a partial observation $o_{t}\sim\mathcal{O}(\cdot|s_{t})$ for current state $s_{t}\in\mathcal{S}$. The agent then selects an action $a_{t}\in\mathcal{A}$ according to $o_{t}$ or its action-observation history $\tau_{t}:\{o_{0},a_{0},o_{1},a_{1}...,o_{t}\}$. The state transitions to the next state $s_{t+1}$ according to $\mathcal{P}(s_{t+1}|s_{t},a_{t})$, and the agent receives a reward $r_{t}$. The goal for the agent is to find the optimal policy $\pi^{*}:\tau\to\Delta(\mathcal{A})$ that maximizes the return, expressed as $\pi^{*}=\text{arg}\max_{\pi}V_{\pi}$, where $V_{\pi}=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|\pi]$ represents cumulative rewards.
Assuming that the state $s$ is available during training, we can train a policy $\mu:s\to\Delta(\mathcal{A})$ based on this state information. For convenience, while the observations available during training could be any form of privileged information, we directly refer to the state $s$ in the remainder of this paper.
@@ 2.1 Teacher-Student Learning
Since we consider both training the teacher and student, in this paper, we use the term Teacher-Student Learning (TSL) to broadly refer to Imitation Learning (IL) (hussein2017imitation, ), policy distillation (pmlr-v89-czarnecki19a, ), and related approaches, as there is no fundamental distinction between them. In TSL, the teacher policy is typically pre-trained using RL or derived from other methods such as a classical controller, which is assumed to effectively accomplish the desired task. The goal is for the teacher to somehow provide supervision to the student in learning to solve the same task.
A straightforward approach to training the agent is to directly supervise the student’s policy $\pi$ using the teacher’s policy $\mu$, similar to Behavioral Cloning (BC) (pomerleau1991efficient, ; torabi2018behavioral, ):
$\min_{\pi}\mathbb{E}_{s\sim d_{\mu}}[D_{\text{KL}}(\mu(\cdot|s),\pi(\cdot|s))],$ (1)
where $d_{\mu}$ is the distribution of states under the teacher’s policy, and $D_{\text{KL}}$ is the Kullback-Leibler (KL) divergence. This objective encourages the student’s policy to mimic the teacher’s policy for the observed states.
However, when the teacher’s policy is based on privileged information, the student can only learn the statistical average of the teacher’s actions for each observation $o$. Specifically, this leads to $\pi(\cdot|o)=\mathbb{E}_{d_{\mu}}[\mu(\cdot|s)|o=f(s)]$, where $f(s)$ denotes the observable function of the state (Warrington2020RobustAL, ; 10.5555/3540261.3541724, ). This limitation can cause sub-optimal performance, as the student is unable to fully replicate the teacher’s behavior in situations where the unobserved state is crucial for making optimal decisions. We will illustrate this limitation with two examples in the next subsection.
@@ 2.2 Didactic Examples
TigerDoor. In the classic TigerDoor problem (10.5555/3091622.3091666, ), there are two doors with a tiger hidden behind one of them. The possible state $s_{L}$ (tiger behind the left door) and $s_{R}$ (tiger behind the right door), with equal probabilities for each, form $\mathcal{S}=\{s_{L},s_{R}\}$. The action set is $\mathcal{A}=\{a_{L},a_{R},a_{l}\}$, where $a_{L}$ and $a_{R}$ denote opening the left and right doors, respectively, and $a_{l}$ denotes listening to determine the tiger’s location. The teacher knows the tiger’s location whereas the student can only ascertain it after choosing $a_{l}$. The payoff matrix is shown in Table 2. The optimal policy for the teacher is to always choose the correct door without listening, whereas the student’s optimal strategy involves first listening to locate the tiger. Consequently, the student cannot learn the optimal policy through supervision from the teacher, as the teacher never chooses $a_{l}$. Under the teacher’s supervision, the student will only learn to randomly select between $a_{L}$ and $a_{R}$, resulting in an expected reward of 0.5. This scenario poses challenges for the supervised student, as the teacher fails to explore and gather essential information for the learner.
@@
TigerDoor-alt. We introduce an alternative version of the problem, called TigerDoor-alt, which also highlights an imitation gap, even without additional exploratory information. In this scenario, the listening action $a_{l}$ is removed, and the reward for correctly selecting the left door is increased to 2 as shown in Table 3. Similarly, the teacher continues to select the correct door, while the student learns to randomly choose between the two doors, yielding an expected reward of 0.75. However, the optimal policy for the student is to always choose the left door, which provides an expected reward of 1. This discrepancy arises from the loss of information when converting the reward-based objective into a policy-supervised objective.
@@
The current solution to this issue is to incorporate environmental rewards into the student’s learning process. To effectively utilize teacher supervision, there are generally two kinds of approaches. The first type dynamically adjusts the weight of the supervision loss between the teacher and pure RL training. This allows the algorithm to switch to pure RL when the teacher is deemed “impossibly good” (10.5555/3540261.3541724, ; pmlr-v202-shenfeld23a, ; shenfeld2023tgrl, ). However, such approaches fail to fully utilize privileged information and may waste the valuable, often expensive, pre-trained teacher. The second kind incorporates teacher supervision into the reward signal, for instance, by using reward shaping via the teacher’s value function (Walsman2023ImpossiblyGE, ). However, this supervision is indirect and may require additional learning. Crucially, none of the existing methods provide a theoretical guarantee that teacher supervision will actually be beneficial. In the following sections, we introduce a method that guarantees the teacher will be a good guide, enabling more effective use of privileged information and supervision to train the student. For a detailed discussion of related work, please refer to Appendix A.
@@ 3 Method
We present our Guided Policy Optimization (GPO) framework, which co-trains two entities: the guider and the learner, which we use to ease the presentation and differentiate from existing TSL methods. Inspired by GPS, GPO iteratively updates both policies to ensure alignment. We then explore both the theoretical properties and practical implementation of GPO, introducing two variants: GPO-penalty and GPO-clip.
@@ 3.1 From GPS to GPO
Unlike direct policy search methods, GPS does not optimize policy parameters directly. Instead, it introduces an intermediate agent (guider) and employs trajectory optimization to learn a time-varying linear-Gaussian policy, which is then used to train a neural network policy (learner) through supervised learning. The GPS procedure has two key phases:
  • Control Phase: A control policy interacts with the environment to minimize costs while ensuring learnability by the neural network policy.
  • Supervised Phase: The neural network policy is trained to mimic the control policy similar to BC.
Although GPS is a model-based method and is not directly applicable in our setting, its core idea of introducing an intermediate agent to guide policy learning can be extended to an RL algorithm in the context of POMDPs.
Specifically, since the guider is only used during training, it can access any type of privileged information. The key requirement is to ensure that the guider is imitable by the learner, which motivates us to introduce the GPO framework, which operates through the following four key steps:
@@
Figure 1: The comparison between TSL and GPO.
The comparison between TSL and GPO is illustrated in Fig. 1. Several key differences between the two frameworks exist. First, the teacher in TSL is typically provided or trained independently from the student, while in GPO, the guider and learner are trained together. Second, TSL typically allows the student to interact with the environment, whereas GPO only uses a guider, enabling more effective trajectory collection due to the behavioral policy being conditioned on privileged information. Lastly, and most importantly, TSL does not use the student to constrain the teacher. This means that if the teacher is too advanced for the student, the student will struggle to learn from the teacher.
@@
$\hat{\mu}=\text{arg}\min\{-\eta_{k}\langle\nabla V(\mu^{(k)}),\mu\rangle+D_{\mu^{(k)}}(\mu,\mu^{(k)})\},$
@@
$\pi^{(k+1)}=\underset{\pi\in\Pi}{\text{arg}\min}\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\pi\rangle+D_{\pi^{(k)}}(\pi,\pi^{(k)})\}.$
@@
Here, we assume that the guider $\mu$ has access to an unlimited policy class, while the learner $\pi$ is constrained to a limited policy class $\Pi$ for simplicity. Policy mirror descent (DBLP:journals/corr/abs-2005-09814, ; JMLR:v23:22-0056, ) is a general family of algorithms that encompasses a wide range of fundamental methods in RL, including trust-region algorithms like TRPO (DBLP:journals/corr/SchulmanLMJA15, ) and PPO (DBLP:journals/corr/SchulmanWDRK17, ). This proposition shows that, despite the learner not directly interacting with the environment, the GPO update for the learner can be viewed as a standard RL update. Specifically, if we use trust-region RL algorithms for the guider, the update for the learner’s policy inherits the key properties, such as policy improvement (DBLP:journals/corr/SchulmanLMJA15, ). This suggests that GPO can effectively address challenges in TSL, such as dealing with a suboptimal teacher or the imitation gap, while still framing the learner’s policy as being supervised by the guider. In Appendix D, we provide an intuitive example illustrating how GPO can achieve optimal in the TigherDoor-alt problem.
Given that GPO mirrors direct RL for the learner, one may ask: What are GPO’s key advantages? The main benefit lies in leveraging additional information while simplifying learning. Since policy gradients suffer from high variance—especially under partial observability—GPO splits learning into two phases: the guider with privileged information handles complex RL via policy gradients, while the partial observable learner is trained via an easier supervised learning, reducing variance and complexity. For instance, to train robustness to noisy observations, GPO can train the guider on clean inputs and supervise the learner with noisy ones, resulting in a more stable and effective learning process.

### v2

@@ 1 Introduction
Many real-world tasks can be formulated as sequential decision-making problems where agents take actions in an environment to achieve specific goals over time (Puterman, 2014). Reinforcement Learning (RL) has emerged as a powerful tool for solving such tasks, leveraging trial-and-error learning to optimize long-term rewards (Sutton & Barto, 2018). Despite its success, RL encounters significant hurdles in complex and partially observable environments, where agents often operate with limited or noisy information (Madani et al., 1999). However, during training, we often have access to extra information that could significantly enhance learning efficiency and performance (Lee et al., 2020a; Chen et al., 2022). For instance, in robotics, while real-world sensor data may be noisy or incomplete, simulation environments typically provide full state observability.
@@
Despite the potential of such privileged information, effectively leveraging it in practice remains a major challenge. One popular strategy to utilize this information is through methods like Imitation Learning (IL) (Hussein et al., 2017), Teacher-Student Learning (TSL), or policy distillation (Czarnecki et al., 2019). In these approaches, a teacher, equipped with privileged information, provides supervision to guide the student’s learning process. However, this strategy introduces its own set of challenges: a teacher with privileged information may impose an unrealistically high-performance standard, making it difficult for the student to effectively imitate. This issue, known as the “impossibly good” teacher (Walsman et al., 2023) or imitation gap (Weihs et al., 2024), can hinder learning and degrade performance. To address this, previous work has sought to integrate environmental rewards into the learning process of the student. One approach is to combine RL with IL (Weihs et al., 2024; Shenfeld et al., 2023a; Nguyen et al., 2023), switching to RL-based training when the teacher becomes inimitable. Another approach modifies environmental rewards based on the teacher through policy distillation (Czarnecki et al., 2019; Walsman et al., 2023). However, such methods diminish the utility of privileged information, often resulting in inefficient use of the teacher’s knowledge.
@@
To better exploit available information, we propose training a “possibly good" teacher, i.e., a teacher whose policy remains within the learner’s imitable region. Inspired by Guided Policy Search (GPS) (Levine & Koltun, 2013; Montgomery & Levine, 2016), we introduce Guided Policy Optimization (GPO), a novel framework that trains both the teacher and the student simultaneously while ensuring that the teacher’s policy remains aligned with that of the student. The key insight behind GPO is that by leveraging privileged information during training, the teacher can be trained more effectively while ensuring that its performance is “possibly good," thus facilitating easier imitation by the student. Theoretically, we show that the student can achieve optimality similar to direct RL training, mitigating the suboptimality and imitation gaps that often arise from purely teacher-based supervision. We empirically validate our algorithm across various tasks, including didactic examples, challenging continuous control tasks in partially observable, noisy environments within the Brax (Freeman et al., 2021) domain, and in memory-based tasks from the POPGym (Morad et al., 2023) benchmark. GPO shows consistent and significant improvements, underscoring its ability to exploit extra information and deliver robust performance across diverse domains.
@@ 2 Background
We consider Partially Observable Markov Decision Process (POMDP) (Kaelbling et al., 1998), which is characterized by the tuple $\langle\mathcal{S},\mathcal{A},r,\mathcal{P},\mathcal{O},\gamma\rangle$. $\mathcal{S}$ represents the set of states, $\mathcal{A}$ the set of actions, $r$ the reward function, $\mathcal{P}$ the transition probability function, $\mathcal{O}$ the partial observation function and $\gamma$ the discount factor. At each time step $t$, the agent receives a partial observation $o_{t}\sim\mathcal{O}(\cdot|s_{t})$ for current state $s_{t}\in\mathcal{S}$. The agent then selects an action $a_{t}\in\mathcal{A}$ according to $o_{t}$ or its action-observation history $\tau_{t}:\{o_{0},a_{0},o_{1},a_{1}...,o_{t}\}$. The state transitions to the next state $s_{t+1}$ according to $\mathcal{P}(s_{t+1}|s_{t},a_{t})$, and the agent receives a reward $r_{t}$. The goal for the agent is to find the optimal policy $\pi^{*}:\tau\to\Delta(\mathcal{A})$ that maximizes the return, expressed as $\pi^{*}=\text{arg}\max_{\pi}V_{\pi}$, where $V_{\pi}=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|\pi]$ represents cumulative rewards. When full state information $s$ is available during training, we may also define a policy $\mu:s\to\Delta(\mathcal{A})$ based on privileged information. For clarity, throughout this paper we refer to such privileged training inputs simply as the state $s$, though in practice they could take other forms. Likewise, we refer to partial observations simply as $o$, though in practice they may include histories or other derived features.
Finally, we emphasize that in the remainder of this paper, the term “optimal” refers to the student’s optimal policy under partial observability—not the teacher’s optimal policy under privileged information, which is generally unattainable for the student.
@@ 2.1 Teacher-Student Learning
Since we consider both training the teacher and student, in this paper, we use the term Teacher-Student Learning (TSL) to broadly refer to Imitation Learning (IL) (Hussein et al., 2017), policy distillation (Czarnecki et al., 2019), and related approaches, as there is no fundamental distinction between them. In TSL, the teacher policy is typically pre-trained using RL or derived from other methods such as a classical controller, which is assumed to effectively accomplish the desired task. The goal is for the teacher to somehow provide supervision to the student in learning to solve the same task.
A straightforward approach to training the agent is to directly supervise the student’s policy $\pi$ using the teacher’s policy $\mu$, similar to Behavioral Cloning (BC) (Pomerleau, 1991; Torabi et al., 2018):
$\min_{\pi}\mathbb{E}_{s\sim d_{\mu}}[\text{D}_{\text{KL}}(\mu(\cdot|s),\pi(\cdot|s))],$ (1)
where $d_{\mu}$ is the distribution of states under the teacher’s policy, and $\text{D}_{\text{KL}}$ is the Kullback-Leibler (KL) divergence. This objective encourages the student’s policy to mimic the teacher’s policy for the observed states. However, when the teacher’s policy is based on privileged information, the student can only learn the statistical average of the teacher’s actions (Warrington et al., 2020; Weihs et al., 2024), and be strictly suboptimal (Cai et al., 2024). In this paper, we refer to such a teacher as inimitable, and we highlight this limitation through two illustrative examples in the next subsection.
@@ 2.2 Didactic Examples
state action $a_{L}$ $a_{R}$ $a_{l}$
$s_{L}$ 1 0 -0.1
$s_{R}$ 0 1 -0.1
Table 1: TigerDoor problem
state action $a_{L}$ $a_{R}$
$s_{L}$ 2 0
$s_{R}$ 0 1
Table 2: TigerDoor-alt problem
@@
TigerDoor. In the classic TigerDoor problem (Littman et al., 1995), there are two doors with a tiger hidden behind one of them. The possible state $s_{L}$ (tiger behind the left door) and $s_{R}$ (tiger behind the right door), with equal probabilities for each, form $\mathcal{S}=\{s_{L},s_{R}\}$. The action set is $\mathcal{A}=\{a_{L},a_{R},a_{l}\}$, where $a_{L}$ and $a_{R}$ denote opening the left and right doors, respectively, and $a_{l}$ denotes listening to determine the tiger’s location. The teacher knows the tiger’s location whereas the student can only ascertain it after choosing $a_{l}$. The payoff matrix is shown in Table 2. The optimal policy for the teacher is to always choose the correct door without listening, whereas the student’s optimal strategy involves first listening to locate the tiger. Consequently, the student cannot learn the optimal policy through supervision from the teacher, as the teacher never chooses $a_{l}$. Under the teacher’s supervision, the student will only learn to randomly select between $a_{L}$ and $a_{R}$, resulting in an expected reward of 0.5. This scenario poses challenges for the supervised student, as the teacher fails to explore and gather essential information for the learner.
@@
TigerDoor-alt. We also introduce an alternative version of the problem, TigerDoor-alt (Table 2), which further illustrates the imitation gap, even when no exploratory actions are required. In this scenario, the listening action $a_{l}$ is removed, and the reward for correctly selecting the left door is increased to 2. Similarly, the teacher continues to select the correct door, while the student learns to randomly choose between the two doors, yielding an expected reward of 0.75. However, the optimal policy for the student is to always choose the left door, which provides an expected reward of 1. This discrepancy arises from the loss of information when converting the reward-based objective into a policy-supervised objective.
@@
The current solution to this issue is to incorporate rewards into the student’s learning process. To effectively utilize teacher supervision, there are two kinds of approaches. The first type dynamically adjusts the weight of the supervision loss between the teacher and pure RL training. This allows the algorithm to switch to pure RL when the teacher is deemed inimitable (Weihs et al., 2024; Shenfeld et al., 2023a; b). However, such approaches fail to fully utilize privileged information and may waste the valuable, often expensive, pre-trained teacher. The second kind incorporates teacher supervision into the reward signal, for instance, by using reward shaping via the teacher’s value function (Walsman et al., 2023). However, this supervision is indirect and may require additional learning. Crucially, to the best of our knowledge, none of the existing methods provide theoretical guarantees that teacher supervision will actually be beneficial.
Another research direction attempts to reconstruct privileged information from partial observations. However, such methods require the MDP to be decodable (Efroni et al., 2022), which is clearly infeasible in the TigerDoor setting. For a more detailed discussion of related work, see Appendix A.
@@ 3 Method
We present our Guided Policy Optimization (GPO) framework, which co-trains two entities: the guider and the learner, which we use to differentiate from existing TSL methods. GPO iteratively updates both policies to ensure alignment. We then explore both the theoretical properties and practical implementation of GPO, introducing two variants: GPO-penalty and GPO-clip.
@@ 3.1 From GPS to GPO
Unlike direct policy search methods, GPS (Levine & Koltun, 2013; Montgomery & Levine, 2016) does not optimize policy parameters directly. Instead, it introduces an intermediate agent (guider) and employs trajectory optimization to learn a time-varying linear-Gaussian policy, which is then used to train a neural network policy (learner) through supervised learning. Although GPS is a model-based method and is not directly applicable in our setting, its idea of introducing an intermediate agent to guide policy learning can be extended to an RL algorithm in the context of POMDPs.
Specifically, since the guider is only used during training, it can access any type of privileged information. The key requirement is to ensure that the guider is imitable by the learner, which motivates us to introduce the GPO framework, which operates through the following four steps:
@@
The comparison between TSL and GPO is illustrated in Fig. 1. Several differences between the two frameworks exist. First, the teacher in TSL is typically provided or trained independently from the student, while in GPO, the guider and learner are trained together. Second, TSL typically allows the student to interact with the environment, whereas GPO only uses a guider, enabling more effective trajectory collection due to the behavioral policy being conditioned on privileged information. Lastly, and most importantly, TSL does not use the student to constrain the teacher. This means that if the teacher is too advanced for the student, the student will struggle to learn from the teacher.
@@
$\hat{\mu}=\text{arg}\min\{-\eta_{k}\langle\nabla V(\mu^{(k)}),\mu\rangle+\text{D}_{\mu^{(k)}}(\mu,\mu^{(k)})\},$
@@
$\pi^{(k+1)}=\underset{\pi\in\Pi}{\text{arg}\min}\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\pi\rangle+\text{D}_{\pi^{(k)}}(\pi,\pi^{(k)})\}.$
@@
Here, we assume that the guider $\mu$ has access to an unlimited policy class, while the learner $\pi$ is constrained to a limited policy class $\Pi$ for simplicity. Policy mirror descent (Tomar et al., 2020; Xiao, 2022) is a general family of algorithms that encompasses a wide range of fundamental methods in RL, including trust-region algorithms like TRPO (Schulman et al., 2015a) and PPO (Schulman et al., 2017). This proposition shows that, despite the learner not directly interacting with the environment, the GPO update for the learner can be viewed as a standard RL update. Specifically, if we use trust-region RL algorithms for the guider, the update for the learner’s policy inherits the key properties, such as policy improvement (Schulman et al., 2015a). This suggests that GPO can effectively address challenges in TSL, such as dealing with a suboptimal teacher or the imitation gap, while still framing the learner’s policy as being supervised by the guider. In Appendix D, we provide an intuitive example illustrating how GPO can achieve optimal in the TigerDoor-alt problem.
Given that GPO mirrors direct RL for the learner, one may ask: What are GPO’s key advantages? The main benefit lies in leveraging additional information while simplifying learning. Since policy gradients suffer from high variance especially under partial observability, GPO splits learning into two phases: the guider with privileged information handles complex RL gradients, while the partial observable learner is trained via an easier supervised learning, reducing variance and complexity. For instance, to train robustness to noisy observations, GPO can train the guider on clean inputs and supervise the learner with noisy ones, resulting in a more stable and effective learning process.